A car uses 8 gallons of gasoline on a trip of 80 miles. At that rate, how much gasoline will a trip of 160 miles require?
Answer by checkley71(8403)

80/8=160/x
80x=8*160
80x=1280
x=1280/80
x=16 gallons of gas would be required for a trip of 160 miles.